OpenStudy (anonymous):

Say we have a diameter of 22.1 mm with uncertainty +-0.1mm Now we have to calculate the percentage uncertainty in area.

OpenStudy (anonymous):

My question is basically what happens to the uncertainty when you divide the diamater by 2 to get the radius?

OpenStudy (mokeira):

what is the figure of the object first?

OpenStudy (anonymous):

A circle

OpenStudy (mokeira):

basically this is what I would do, I hope you understand First I would find he area of the circle let area=x then I would say y=x-0.1 z=x+0.1 then add the uncertainties... y+z after adding them up, divide them by the area and multiply by 100...and you have the percentage uncertainty! :)

OpenStudy (anonymous):

What I am asking is that when you divide diameter 22.1mm by 2, if anything happens to the uncertainty?

OpenStudy (mokeira):

are you asking if uncertainty will double or something like that?

OpenStudy (mokeira):

no it would be constant

OpenStudy (mokeira):

i think

OpenStudy (anonymous):

So if it's a constant then nothing happens to it?

OpenStudy (anonymous):

Nothing happens to the percentage uncertainty

OpenStudy (mokeira):

the diameter will have the same uncertainty as the full circle but you can try it practically find are of a circle then find it with half the measurement and see

OpenStudy (anonymous):

Ok I see. Thanks for your help!